OpenStudy (anonymous):

How do I set up this problem? "If a ball is thrown straight up into the air with an initial velocity of 45 ft/s, its height in feet after t seconds is given by y=45t-16t^2. Find the average velocity for the time period beginning when t=1 and lasting 0.5 seconds."

OpenStudy (aaronq):

The average velocity is the difference in position over the time interval: \(\large v_{avg}=\dfrac{y(t)_f-y(t)_i}{\Delta t}=\dfrac{y(1)-y(0.5)}{1-0.5~s}\) \(\large \large v_{avg}=\dfrac{[45(1)-16(1)^2]m-[45(0.5)-16(0.5)^2]m}{1-0.5~s}\)